OpenStudy (anonymous):

Sound source of frequency 170 Hz is placed near a wall. A man walks from the source to the wall and finds periodic rise and fall in sound intensity.Speed of sound in air is 340m/s. Distance separating two adjacent minimum intensity positions in metres is? I get the (correct)answer 1 if I consider standing waves but I want to know if and how beats will be produced here and if that concept is used.

OpenStudy (lastdaywork):

I think beats and standing waves will give the same answer. But to be sure, I need to revise my notes (and I am wasted right now). @Mashy ; @Vincent-Lyon.Fr ; @douglaswinslowcooper

OpenStudy (anonymous):

\[\phi(x,t)=j(vector)\sin(2\pi vt/\lambda) \cos(2\pi x/\lambda) \]

OpenStudy (anonymous):

 represents a transverse standing wave how?